# Adaptive Width Neural Networks

### v2

@@ 6 Conclusions
We introduced a new methodology to learn an unbounded width of neural network layers within a single training, by imposing a soft ordering of importance among neurons. Our approach requires very few changes to the architecture, adapts the width to the task’s difficulty, and does not impact negatively convergence, although it introduces some overhead during training. We showed stability convergence to similar widths under bounded activations for different hyper-parameters configurations, advocating for a practical reduction of the width’s search space. A by-product of neurons’ ordering is the ability to easily compress the network during or after training, which is relevant in the context of foundational models trained on large data, which are believed to require billions of parameters. Finally, we have tested AWNN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved the same loss with 200x less parameters.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Full derivation of the rescaled weight initialization
@@
Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in Glorot & Bengio (2010)
@@
Finally, because the mean of the independent variables involved is zero by assumption, it holds that $\mathrm{Var}[w^{\ell}_{ij}\alpha^{\ell-1}_{j}]=\mathrm{Var}[w^{\ell}_{ij}]\mathrm{Var}[\alpha^{\ell-1}_{j}]$. We can also abstract from the indexes, since the weight variables are i.i.d. and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^3] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of Glorot & Bengio (2010), since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$. In Glorot & Bengio (2010) our $D_{\ell-1}$ is denoted as “$n_{i^{\prime}}$” (see Equation 5).
@@
From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^4] Note that what we impose is different from Glorot & Bengio (2010), where the specific position $i$ is irrelevant. Here, we are asking that the variance of the gradients for neurons in the same position $i$, but at different layers, stays constant. Alternatively, we could impose an equivalence for all $i\neq i^{\prime}$..
Using Equation 18, and considering as in Glorot & Bengio (2010) that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
@@
$\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right]=(p^{\ell}_{i})^{2}\left(\prod_{i=\ell+1}^{\ell+2}\mathrm{Var}[w^{\ell+1}]\right)\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\underbrace{\left(\sum_{j^{\prime}=1}^{D_{\ell+2}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{\ell+2})}{\partial\alpha_{j^{\prime}}^{\ell+2}}\right]\right)}_{\text{constant w.r.t.\ }j}$ (32)
@@
Both Equations 28 and 37 are verified when we initialize the distributions $f_{\ell}$ in the same way for all layers, which implies $(p^{\ell}_{i})^{2}=(p^{\ell+1}_{i})^{2}$. Note that without this last requirement, Equation 37 would violate the i.i.d. assumption of the weights.
@@
Using Lemma A.1, we can rewrite Equation 22 as
$\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]=\frac{1}{2}\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]=\frac{1}{2}\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ (40)
@@
From a backpropagation perspective, since $\sigma^{\prime}(x)\not\approx 1$, we use similar arguments as in He et al. (2015): we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability. One can show, by expanding the definition of variance, that $\mathrm{Var}[\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}]=\frac{1}{2}\mathrm{Var}[w^{\ell+1}_{ji}]$. As a result, one can re-write Equation 22 as
@@
where we have made a rough assumption of independence between $\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)$ and $w^{\ell+1}_{ji}$ since the former term is likely a constant. Following an identical derivation as above, we obtain
@@
Below, we provide a comparison of convergence between an AWNN ReLU MLP initialized with the standard kaiming scheme and one initialized according to our theoretical results (“kaiming+”).
@@
Appendix B Dataset Info and Statistics
@@
Note that for Multi30k, the maximum sequence length is 128 and the number of possible tokens is 50527. The classification loss is computed by comparing the predicted token against the expected one, ignoring any token beyond the length of the actual target. For a detailed description of the graph datasets, the reader can refer to Errica et al. (2020).
Appendix C Hyper-parameters for fixed and AWNN models
For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWNN. We try our best to keep architectural choices identical with the exception of the width of the MLPs used by each model. In all experiments, we either run patience-based early stopping (Prechelt, 1998) or simply select the epoch with best validation score.
C.1 Synthetic Tabular Datasets
We report the hyper-parameter configurations tried for the fixed MLP and AWNN in Table 3. In particular, the starting width of AWNN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons. We also noticed that a learning rate value of $0.1$ makes learning the width unstable, so we did not use it in our experiments. The total number of configurations for the fixed MLP is 180, for AWNN is 18. We also did not impose any prior on an expected size to see if AWNN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
Table 3: Hyper-parameter configurations for standard MLP and AWNN versions on tabular datasets.

### v3

@@ 6 Conclusions
We introduced a new methodology to learn an unbounded width of neural network layers within a single training, by imposing a soft ordering of importance among neurons. Our approach requires very few changes to the architecture, adapts the width to the task’s difficulty, and does not impact negatively convergence. We showed stability of convergence to similar widths under bounded activations for different hyper-parameters configurations, advocating for a practical reduction of the width’s search space. A by-product of neurons’ ordering is the ability to easily compress the network during or after training, which is relevant in the context of foundational models trained on large data, which are believed to require billions of parameters. Finally, we have tested AWNN on different models and data domains to prove its broad scope of applicability: a Transformer architecture achieved a similar loss with 200x less parameters.
Appendix A Background Notions
Since we were mostly inspired by the work of  Nazaret and Blei [42], we complement the main text with an introduction to truncated distributions. In particular, by truncating a distribution at its quantile function evaluated at $k$ (see Appendix B for a visual explanation), the support of the distribution becomes finite and countable, hence we can compute expectations in a finite number of steps. A truncated distribution should ideally satisfy some requirements defined in  Nazaret and Blei [42] and illustrated below.
Definition A.1.
A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected and bounded members if
  1. 1.
    $\forall q\in Q$, $\operatorname{support}(q)$ is bounded
  2. 2.
    $\forall L\in\mathbb{N}^{+},\exists q\in Q$ such that $L\in\operatorname{argmax}(q)$
  3. 3.
    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
The first condition allows us to compute the expectation over any $q\in Q$ in finite time, condition 2 ensures there is a parametrization that assigns enough probability mass to each point in the support of $q$, and condition 3 is necessary for learning via backpropagation.
An important distinction with  Nazaret and Blei [42] is that we do not need nor want condition 2 to be satisfied. As a matter of fact, we want to enforce a strict ordering of neurons where the first is always the most important one. Condition 2 is useful, as done in  Nazaret and Blei [42], to assign enough ‘importance” to a specific layer of an adaptive-depth architecture, while previous layers are still functional to the final result. In the context of adaptive-width networks, however, this mechanism may leave a lot of neurons completely unutilized, letting the network grow indefinitely. That is why we require that each distribution $q$ is monotonically decreasing and all possible widths should have non-zero importance. We change the definition for the distributions we are interested in as follows:
Definition A.2.
A variational family $Q={q(\boldsymbol{\omega})}$ over $\mathbb{N}^{+}$ is unbounded with connected, bounded, and decreasing members if
  1. 1.
    $\forall q\in Q$, $\operatorname{support}(q)$ is bounded
  2. 2.
    $\forall q\in Q$, $q$ is monotonically decreasing and $\forall x\in\mathbb{N}^{+},q(x)>0$
  3. 3.
    Each parameter in the set $\boldsymbol{\omega}$ is a continuous variable.
Appendix B How to Compute $D_{\ell}$ in Practice
To aid the understanding of how one computes $D_{\ell}$, we provide a visualization of the exponential distribution, its discretized version (Equation 6) and the quantile function evaluated at $k=0.9$ in Figure 6 below.
Figure 6: On the left, we provide a visualization of the exponential density with $\lambda=0.5$, its cumulative density, and the quantile function evaluated at $k=0.9$, which returns a value close to $5$. On the right, we show the discretized exponential and its cumulative mass function; we can see that the quantile function of the continuous density provides an upper bound for the natural where the cumulative mass function is greater than $k=0.9$. Therefore, one can safety use the ceiling of the output of the quantile function to determine the width $D_{\ell}$ to use for a given value of $\lambda$.
Appendix C Full ELBO Derivation
The full ELBO derivation is as follows:
$\mathbb{E}_{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\left[\log\frac{p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)}{q(\boldsymbol{\lambda},\boldsymbol{\theta})}\right]=$ (11)
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}[\log p(Y,\boldsymbol{\lambda},\boldsymbol{\theta}|X)-\log q(\boldsymbol{\lambda},\boldsymbol{\theta})]=$ (12)
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log(p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)p(\boldsymbol{\theta})p(\boldsymbol{\lambda}))-\log(q(\boldsymbol{\lambda})q(\boldsymbol{\theta}|\boldsymbol{\lambda}))]$ (13)
where we factorized probabilities according to the independence assumptions of the graphical model. Then
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)+\log p(\boldsymbol{\lambda})+\log p(\boldsymbol{\theta})-\log q(\boldsymbol{\lambda})-\log q(\boldsymbol{\theta}|\boldsymbol{\lambda})]=$ (14)
$\mathbb{E}_{q(\boldsymbol{\lambda})}\mathbb{E}_{q(\boldsymbol{\boldsymbol{\theta}|\lambda})}[\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)]\approx$ (15)
$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)]$ (16)
where we use the first-order approximation in the last step. Note that when computing $\frac{p(\theta)}{q(\theta|\lambda)}$, the products from $D_{\ell+1}$ to infinity cancel out. Equation 13 follows by simply applying definitions, the iid assumption $p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)=\prod_{i}^{N}p(y_{i}|\boldsymbol{\nu},\boldsymbol{\rho},x_{i})$, and by transforming products into summations using the logarithm.
An important note about the first-order (Taylor) approximation $E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})q(\boldsymbol{\theta}|\boldsymbol{\lambda};\boldsymbol{\rho})}[f(\boldsymbol{\lambda},\boldsymbol{\theta})]\approx E_{q(\boldsymbol{\lambda};\boldsymbol{\nu})}[f(\boldsymbol{\lambda},\boldsymbol{\rho})]\approx f(\boldsymbol{\nu},\boldsymbol{\rho})$ is that the function $f(\boldsymbol{\lambda},\boldsymbol{\theta})=\log\frac{p(\boldsymbol{\lambda})}{q(\boldsymbol{\lambda})}+\log\frac{p(\boldsymbol{\theta})}{q(\boldsymbol{\theta}|\boldsymbol{\lambda})}+\log p(Y|\boldsymbol{\lambda},\boldsymbol{\theta},X)$ is not differentiable with respect to $D_{\ell}$ when there is a jump in the quantile evaluated at $k$, so it would seem we cannot apply it. This also happens in  Nazaret and Blei [42], Equation 11, when the function is a neural network containing non-differentiable activations like ReLU, but we could not find a discussion about it. Importantly, the function is not differentiable at a countable set of points (one for each possible width), which has a measure null. Since the Lebesgue integral of the expectation is over reals, we should be able to remove this set from the integral while ensuring that the approximation still holds mathematically.
Appendix D Lipschitz Continuity of the ELBO
Whenever the quantile of the finite distribution changes, we need to change the number of neurons of the neural network. Here, we want to show that the change in the ELBO after a change in the number of neurons is bounded. Therefore, we provide a theoretical result (Proposition D.1) that shows that the ELBO satisfies the Lipschitz continuity.
Proposition D.1 (AWNN ELBO Lipschitz continuity).
The ELBO loss of Equation 9, with respect to the change in the depth $D_{\ell}$ for the layer $l$, is Lipschitz continuous.
Proof.
We focus on the term involving $D_{\ell}$ of the ELBO, we write Equation 9 as
$\log\frac{p(\boldsymbol{\nu})}{q(\boldsymbol{\nu})}+\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}+\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
where only the second and last terms depend on $\mathbf{D}=\{D_{\ell}\}_{\ell=1}^{L}$ Let’s first define
$\log\frac{p(\boldsymbol{\rho})}{q(\boldsymbol{\rho}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}\sum_{n=1}^{D_{\ell}}\log\frac{p(\rho^{\ell}_{n})}{q(\rho^{\ell}_{n}|\boldsymbol{\nu})}=\sum_{{\ell}=1}^{L}f_{1}(D_{\ell})$
We have that $f_{1}(D_{\ell})$ is Lipschitz continuous, indeed, when $D_{\ell}$ changes to $D_{l^{\prime}}$, we have
$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|$ $=|\sum_{n=D_{\ell}}^{D_{\ell^{\prime}}}\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (17)
$\leq~\sum_{n=D_{l}}^{D_{\ell^{\prime}}}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$ (18)
$\leq\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}||D_{\ell^{\prime}}-D_{\ell}|$ (19)
Therefore
$|f_{1}(D_{\ell^{\prime}})-f_{1}(D_{\ell})|\leq M|D_{\ell^{\prime}}-D_{\ell}|$
with $M=\max_{n}|\log\frac{p(\boldsymbol{\rho}_{n})}{q(\boldsymbol{\rho}_{n}|\boldsymbol{\nu})}|$. We now look at the last term,
$f_{2}(\mathbf{D})=\log p(Y|\boldsymbol{\nu},\boldsymbol{\rho},X)$
This function is a neural network followed by the squared norm. Therefore, $f_{2}(\mathbf{D})$ is continuous almost everywhere  [59], therefore Lipschitz on both parameters and the input. When we change $D_{\ell}$ to $D_{\ell^{\prime}}$, we are adding network parameters, in particular, if we look at the two-layer network, we have
$y=V\sigma(Wx+b)+c,V\in\mathbb{R}^{m\times n},W\in\mathbb{R}^{n\times d}$
we then have new parameters, which can be shown to be
$y^{\prime}=V\sigma(Wx+b)+c+\underbrace{V^{\prime}\sigma(W^{\prime}x+b^{\prime})+c^{\prime}}_{\text{new neurons}},V^{\prime}\in\mathbb{R}^{m\times n^{\prime}},W^{\prime}\in\mathbb{R}^{n^{\prime}\times d},b^{\prime}\in\mathbb{R}^{n^{\prime}},c^{\prime}\in\mathbb{R}^{m}$
the second term is also Lipschitz, therefore the function $f_{2}(\mathbf{D})$ is also Lipschitz. Indeed the network is composed of a linear operation and the element-wise activation functions. If the activation functions are continuous and of bounded gradient, then the whole network is Lipschitz continuous. ∎
Appendix E Full derivation of the rescaled weight initialization
@@
Let us start from the first case of $\sigma^{\prime}(0)\approx 1$. Using the Taylor expansions for the moments of functions of random variables as in  Glorot and Bengio [20]
@@
Finally, because the mean of the independent variables involved is zero by assumption, it holds that $\mathrm{Var}[w^{\ell}_{ij}\alpha^{\ell-1}_{j}]=\mathrm{Var}[w^{\ell}_{ij}]\mathrm{Var}[\alpha^{\ell-1}_{j}]$. We can also abstract from the indexes, since the weight variables are i.i.d. and from that it follows that $\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}[\alpha^{\ell}_{j}]\ \forall i,j,\ \ i,j\in\{1,\dots,D_{\ell}\}$, obtaining [^4] It is worth noting that, in standard MLPs, $p_{j}^{\ell}=1$ so we recover the derivation of  Glorot and Bengio [20], since $\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ would be equal to $D_{\ell-1}$. In  Glorot and Bengio [20] our $D_{\ell-1}$ is denoted as “$n_{i^{\prime}}$” (see Equation 5).
@@
From a backpropagation perspective, a similar desideratum would be to ensure that $\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell})}{\partial\alpha_{i}^{\ell}}\right]=\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{i}^{\ell+1})}{\partial\alpha_{i}^{\ell+1}}\right]$ [^5] Note that what we impose is different from  Glorot and Bengio [20], where the specific position $i$ is irrelevant. Here, we are asking that the variance of the gradients for neurons in the same position $i$, but at different layers, stays constant. Alternatively, we could impose an equivalence for all $i\neq i^{\prime}$..
Using Equation 22, and considering as in  Glorot and Bengio [20] that at initialization we are in a linear regime where $\sigma^{\prime}(x)\approx 1$,
@@
$\mathrm{Var}[w^{\ell+1}](p^{\ell}_{i})^{2}\sum_{j=1}^{D_{\ell+1}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j}^{\ell+1})}{\partial\alpha_{j}^{\ell+1}}\right]$ (36)
$=(p^{\ell}_{i})^{2}\left(\prod_{i=\ell+1}^{\ell+2}\mathrm{Var}[w^{\ell+1}]\right)\sum_{j=1}^{D_{\ell+1}}(p^{\ell+1}_{j})^{2}\underbrace{\left(\sum_{j^{\prime}=1}^{D_{\ell+2}}\mathrm{Var}\left[\frac{\partial\mathcal{L}(\alpha_{j^{\prime}}^{\ell+2})}{\partial\alpha_{j^{\prime}}^{\ell+2}}\right]\right)}_{\text{constant w.r.t.\ }j}$ (37)
@@
Both Equations 32 and 42 are verified when we initialize the distributions $f_{\ell}$ in the same way for all layers, which implies $(p^{\ell}_{i})^{2}=(p^{\ell+1}_{i})^{2}$. Note that without this last requirement, Equation 42 would violate the i.i.d. assumption of the weights.
@@
Using Lemma E.1, we can rewrite Equation 26 as
$\mathrm{Var}[\alpha^{\ell}_{i}]=\mathrm{Var}\left[\sigma\left(\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right)\right]=\frac{1}{2}\mathrm{Var}\left[\sum_{j=1}^{D_{\ell-1}}w^{\ell}_{ij}\alpha^{\ell-1}_{j}p^{\ell-1}_{j}\right]$ (45)
$=\frac{1}{2}\mathrm{Var}[w^{\ell}]\mathrm{Var}[\alpha^{\ell-1}]\sum_{j=1}^{D_{\ell-1}}\left(p^{\ell-1}_{j}\right)^{2}$ (46)
@@
From a backpropagation perspective, since $\sigma^{\prime}(x)\not\approx 1$, we use similar arguments as in  He et al. [25]: we assume that the pre-activations $\alpha$ have zero mean, then the derivative of the ReLU can have values zero or one with equal probability. One can show, by expanding the definition of variance, that $\mathrm{Var}[\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)w^{\ell+1}_{ji}]=\frac{1}{2}\mathrm{Var}[w^{\ell+1}_{ji}]$. As a result, one can re-write Equation 26 as
@@
where we have made the assumption of independence between $\sigma^{\prime}\left(\sum_{k=1}^{D_{\ell}}w^{\ell+1}_{jk}\alpha^{\ell}_{k}p^{\ell}_{k}\right)$ and $w^{\ell+1}_{ji}$ since the former term is likely a constant. Following an identical derivation as above, we obtain
@@
Below, we provide a comparison of convergence between an AWNN ReLU MLP initialized with the standard Kaiming scheme and one initialized according to our theoretical results (“Kaiming+").
@@
Appendix F Dataset Info and Statistics
@@
Note that for Multi30k, the maximum sequence length is 128 and the number of possible tokens is 50527. The classification loss is computed by comparing the predicted token against the expected one, ignoring any token beyond the length of the actual target. For a detailed description of the graph datasets, the reader can refer to  Errica et al. [16].
Appendix G Hyper-parameters for fixed and AWNN models
For each data domain, we discuss the set of hyper-parameters tried during model selection for the fixed baseline and for AWNN. We try our best to keep architectural choices identical with the exception of the width of the MLPs used by each model. In all experiments, we either run patience-based early stopping  [46] or simply select the epoch with best validation score.
G.1 Synthetic Tabular Datasets and Permuted MNIST
We report the hyper-parameter configurations tried for the fixed MLP/RNN and AWNN in Table 3. In particular, the starting width of AWNN under the quantile function evaluated at $k=0.9$ is approximately $256$ hidden neurons. We also noticed that a learning rate value of $0.1$ makes learning the width unstable, so we did not use it in our experiments. The total number of configurations for the fixed models is 180, for AWNN is 18. We also did not impose any prior on an expected size to see if AWNN recovers a similar width compared to the best fixed model (please refer to Table 1) starting from the largest configuration among the fixed models.
Table 3: Hyper-parameter configurations for standard MLP/RNN and AWNN versions on tabular datasets.